Question 5 A researcher wants to find out the effect of marriage on the self‐esteem of couples. He measured the average self‐esteem scores of couples before marriage and a year after their wedding day. Their average self‐esteem scores are shown below.

Pair

After one year

Before marriage

A

14

8

B

16

4

C

18

10

D

10

10

E

9

9

F

16

9

Is there a significant difference in the self‐esteem scores before marriage and a year after the couples’ wedding day? Using the 5 steps of hypothesis testing, test at α .05, one‐tailed test. If the difference is significant, test the effect size through Cohen’s d and describe the effect size.   

**** I am specifically stuck on showing the work of the cohen d calculation of effect size!
Answer #1

Answer:

Question 5 A researcher wants to find out the effect of marriage on the self‐esteem of couples. He measured the average self‐esteem scores of couples before marriage and a year after their wedding day. Their average self‐esteem scores are shown below.

Pair

After one year

Before marriage

A

14

8

B

16

4

C

18

10

D

10

10

E

9

9

F

16

9

Is there a significant difference in the self‐esteem scores before marriage and a year after the couples’ wedding day? Using the 5 steps of hypothesis testing, test at α .05, one‐tailed test. If the difference is significant, test the effect size through Cohen’s d and describe the effect size.   

**** I am specifically stuck on showing the work of the cohen d calculation of effect size!

Null hypothesis Ho: d= 0 Alternative hypothesis: H: Md > 0

Upper tail test

Difference = after-before

u /Ps Pr!

5.5 4.7223/6

= 2.8529

DF = n-1 =5

Table value of t with 5 DF at 0.05 level =2.015

Rejection Region: Reject Ho if t > 2.0150

Calculated t = 2.8529 , falls in the rejection region

The null hypothesis is rejected.

There is a significant difference in the self‐esteem scores before marriage and a year after the couples’ wedding day.

Cohen d is (divide the mean of the differences by the SD of the differences)

Cohen d =5.5/ 4.7223=1.164686699

=1.16

Since calculated Cohen d =1.16 which is greater than 0.8, the effect size is large.

Paired t Test

Data

Hypothesized Mean Difference

0

Level of significance

0.05

Intermediate Calculations

Sample Size

6

DBar

5.5000

Degrees of Freedom

5

SD

4.7223

Standard Error

1.9279

t Test Statistic

2.8529

Upper-Tail Test

Upper Critical Value

2.0150

p-Value

0.0179

Reject the null hypothesis